# Strategy for identifying Vera C. Rubin Observatory kilonova candidates for targeted gravitational-wave searches

### v1

@@ 6 Conclusions
Upcoming contemporaneous observations by the Vera C. Rubin Observatory and the LIGO, Virgo and KAGRA gravitational-wave detectors present an opportunity to detect both kilonovae and gravitational waves from a binary neutron star merger. We consider the case of a search for subthreshold gravitational waves associated with kilonovae observed by Rubin that occur within the sensitive volume of LIGO-Virgo-KAGRA, in analogy with existing searches for gravitational waves associated with gamma-ray bursts and fast radio bursts (R. Abbott et al., 2021b, 2022, 2023c). This effort is complementary to efforts to follow-up gravitational-wave triggers by the Rubin ToO programme (R. Margutti et al., 2018; I. Andreoni et al., 2022a, 2024).
@@
In order to carry out such a search in practice, a sample of kilonova candidates is required. We perform simulations to estimate the population of kilonovae that will be observed by Rubin within 500 Mpc ($z<0.11$), finding that Rubin will detect around 40 kilonovae per year, or 3 to 5 per month. However, most of these kilonovae will not have any alerts emitted, as only 10% will have a detection above $\mathrm{SNR}>5$ for one epoch and will be made public by brokers such as Fink. This number is reduced to 7% when requiring two alerts. Thus the number of strong kilonova candidates is more like 3 per year, in agreement with previous estimates (e.g., I. Andreoni et al., 2019, 2022a; F. Ragosta et al., 2024).
@@
This work has implications for attempts to perform targeted subthreshold searches for gravitational waves associated with kilonovae using the Rubin LSST public alert stream. Our simulations show that typical kilonova candidates detected by Rubin will be detected 1–2 days following the neutron star merger, requiring several days of gravitational-wave data to be searched. This is significantly longer than the seconds of data that are typically searched for gravitational waves when targeting gamma-ray bursts and fast radio bursts (R. Abbott et al., 2021b, 2022, 2023c). This will either require the optimisation of existing search codes, the development of new analyses or the use of suboptimal methods in order to make such a search computationally feasible.
@@
In addition, we also showed that the sample of kilonvoa candidates detected by Rubin/Fink is heavily contaminated by other variables and transients, such as supernovae and asteroids. We have developed selection criteria designed to suppress this contamination as much as possible. We then estimated the number of contaminants that match our selection criteria using archival data from the Zwicky Transient Facility. We find that even using our strictest selection criteria, there will be around 30 events per month matching our selection criteria that are not kilonovae. Our estimate provides a lower limit for the contamination rate for Rubin/LSST. The true contamination rate for Rubin is currently unknown. Based on the ratio of the number of alerts emitted by ZTF ($\sim 2\times 10^{5}$ per night) and Rubin ($\sim 1\times 10^{7}$ per night), it could potentially be up to a factor of 50 worse than this. The true contamination rate from Rubin should be well measured by the start of the Fifth LIGO/Virgo/Kagra observing run. The severe contamination will also increase the computational cost of a subthreshold gravitational wave search. For kilonovoa candidates with a known distance (e.g., from a host galaxy with a known redshift), a binary neutron star merger origin may be able to be excluded if the gravitational wave searches can produce an exclusion distance greater than the distance of the event (cf. J. Abadie et al., 2012).
@@
In contrast to previous analyses of kilonovae in Rubin (e.g., I. Andreoni et al., 2022a), we find that the most common band-passes an alert will be emitted in is $i$ followed shortly by $r$. For programmes such as ToO we encourage reducing the SNR threshold drastically. Only 21% of kilonovae are detected with $\mathrm{SNR}>3$ and up to 50% with $\mathrm{SNR}>2$.
@@
Using the known sky location of the candidate and the approximately known merger time, a more sensitive (‘targeted’) search for gravitational-wave emission associated with a binary neutron star merger can be performed (L. Z. Kelley et al., 2013). Similar previous targeted searches have shown an improvement of up to 25% in distance over an untargeted, all-sky search, leading to a potential doubling of the detection rate (A. R. Williamson et al., 2014). Since Rubin only surveys $\sim 18,000$ deg2 in the Southern hemisphere, we expect that the increase that can be expected is smaller than for all-sky gamma-ray burst monitors. Further studies will be required to assess the feasibility of our proposal, and determine the improvement that can be realised in practice. Targeted searches for gravitational waves associated with kilonova candidates from Rubin may provide the next multi-messenger observation.
@@
We thank Emille Ishida, Julien Peloton and Marion Pillas for useful comments on this manuscript, as well as Rick Kessler for help with SNANA. This research has made use of NASA’s Astrophysics Data System. Software citation information aggregated using The Software Citation Station (T. Wagg & F. S. Broekgaarden, 2024; T. Wagg et al., 2025). The Vera C. Rubin Observatory is funded by the U.S. National Science Foundation and the U.S. Department of Energy’s Office of Science. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. This work was developed within the Fink community and made use of the Fink community broker resources. Fink is supported by LSST-France and CNRS/IN2P3. This research made use of the cross-match service provided by CDS, Strasbourg and the SIMBAS database and has made use of the International Variable Star Index (VSX) database, operated at AAVSO, Cambridge, Massachusetts, USA. This research used resources of the National Energy Research Scientific Computing Center (NERSC). Notebooks producing the plots in this paper can be found at https://github.com/anaismoller/Rubin_KNe_with_Fink_LVK. This research was supported by grants from the Australian Research Council (ARC). AM is supported by the ARC Discovery Early Career Research Award (DE230100055). $S$ is supported by an ARC Discovery Early Career Research Award (DE220100241). Parts of this research were conducted by the ARC Centre of Excellence for Gravitational Wave Discovery (OzGrav), through project number CE230100016.

### v2

@@ 6 Conclusions
Upcoming contemporaneous observations by the Vera C. Rubin Observatory and the LIGO, Virgo and KAGRA gravitational-wave detectors present an opportunity to detect both kilonovae and gravitational waves from a binary neutron star merger. We consider the case of a search for subthreshold gravitational waves associated with kilonovae observed by Rubin that occur within the sensitive volume of LIGO-Virgo-KAGRA, in analogy with existing searches for gravitational waves associated with gamma-ray bursts and fast radio bursts (Abbott:2021ApJLVKGRBO3a; Abbott:2022ApJLVKGRBO3b; Abbott:2023ApJLVKFRBO3a). This effort is complementary to efforts to follow-up gravitational-wave triggers by the Rubin ToO programme (Margutti:2018arXiv; Andreoni:2022ApJS; Andreoni:2024ToO).
@@
In order to carry out such a search in practice, a sample of kilonova candidates is required. We perform simulations to estimate the population of kilonovae that will be observed by Rubin within 500 Mpc ($z<0.11$), finding that Rubin will detect around 40 kilonovae per year, or 3 to 5 per month. However, most of these kilonovae will not have any alerts emitted, as only 10% will have a detection above $\mathrm{SNR}>5$ for one epoch and will be made public by brokers such as Fink. This number is reduced to 7% when requiring two alerts. Thus the number of strong kilonova candidates is more like 3 per year, in agreement with previous estimates (e.g., Andreoni:2019PASP; Andreoni:2022ApJS; Ragosta:2024ApJ).
@@
This work has implications for attempts to perform targeted subthreshold searches for gravitational waves associated with kilonovae using the Rubin LSST public alert stream. Our simulations show that typical kilonova candidates detected by Rubin will be detected 1–2 days following the neutron star merger, requiring several days of gravitational-wave data to be searched. This is significantly longer than the seconds of data that are typically searched for gravitational waves when targeting gamma-ray bursts and fast radio bursts (Abbott:2021ApJLVKGRBO3a; Abbott:2022ApJLVKGRBO3b; Abbott:2023ApJLVKFRBO3a). This will either require the optimisation of existing search codes, the development of new analyses or the use of suboptimal methods in order to make such a search computationally feasible.
@@
In addition, we also showed that the sample of kilonvoa candidates detected by Rubin/Fink is heavily contaminated by other variables and transients, such as supernovae and asteroids. We have developed selection criteria designed to suppress this contamination as much as possible. We then estimated the number of contaminants that match our selection criteria using archival data from the Zwicky Transient Facility. We find that even using our strictest selection criteria, there will be around 30 events per month matching our selection criteria that are not kilonovae. Our estimate provides a lower limit for the contamination rate for Rubin/LSST. The true contamination rate for Rubin is currently unknown. Based on the ratio of the number of alerts emitted by ZTF ($\sim 2\times 10^{5}$ per night) and Rubin ($\sim 1\times 10^{7}$ per night), it could potentially be up to a factor of 50 worse than this. The true contamination rate from Rubin should be well measured by the start of the Fifth LIGO/Virgo/Kagra observing run. The severe contamination will also increase the computational cost of a subthreshold gravitational wave search. For kilonovoa candidates with a known distance (e.g., from a host galaxy with a known redshift), a binary neutron star merger origin may be able to be excluded if the gravitational wave searches can produce an exclusion distance greater than the distance of the event (cf. Abadie:2012ApJ).
An example of a recent kilonova imposter during O4 is SN2025ulz (identified by ZTF; Stein:2025GCN), which was initially considered a possible counterpart to the gravitational-wave trigger S250818k due to its color and fade rate (e.g., Hall:2025arXivEarly). After approximately five days, the transient began to rebrighten, and was subsequently classified as an unrelated type II supernova (Franz:2025arXiv; Gillanders:2025arXiv; Hall:2025arXivDESI; Kasliwal:2025arXiv; OConnor:2025arXiv; Yang:2025arXiv). This classification was only possible due to substantial follow-up of this transient.
@@
In contrast to previous analyses of kilonovae in Rubin (e.g., Andreoni:2022ApJS), we find that the most common band-passes an alert will be emitted in is $i$ followed shortly by $r$. For programmes such as ToO we encourage reducing the SNR threshold drastically. Only 21% of kilonovae are detected with $\mathrm{SNR}>3$ and up to 50% with $\mathrm{SNR}>2$. Blind searches can also benefit from joining different alert streams such as ZTF, LSST and LS4 in particular if the surveys are observing the same patches of the sky. This would require joint stream filtering and potentially benefit from multi-modal machine learning models to identify the most promising candidates.
@@
Using the known sky location of the candidate and the approximately known merger time, a more sensitive (‘targeted’) search for gravitational-wave emission associated with a binary neutron star merger can be performed (Kelley:2013PRD). Similar previous targeted searches have shown an improvement of up to 25% in distance over an untargeted, all-sky search, leading to a potential doubling of the detection rate (Williamson:2014PRDPyGRB). Since Rubin only surveys $\sim 18,000$ deg2 in the Southern hemisphere, we expect that the increase that can be expected is smaller than for all-sky gamma-ray burst monitors. Further studies will be required to assess the feasibility of our proposal, and determine the improvement that can be realised in practice. Targeted searches for gravitational waves associated with kilonova candidates from Rubin may provide the next multi-messenger observation.
@@
We thank Emille Ishida, Julien Peloton and Marion Pillas for useful comments on this manuscript, as well as Rick Kessler for help with SNANA. We thank the annonymous referee for comments that improved the manuscript. This research has made use of NASA’s Astrophysics Data System. Software citation information aggregated using The Software Citation Station (software-citation-station-paper; software-citation-station-zenodo). The Vera C. Rubin Observatory is funded by the U.S. National Science Foundation and the U.S. Department of Energy’s Office of Science. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. This work was developed within the Fink community and made use of the Fink community broker resources. Fink is supported by LSST-France and CNRS/IN2P3. This research made use of the cross-match service provided by CDS, Strasbourg and the SIMBAS database and has made use of the International Variable Star Index (VSX) database, operated at AAVSO, Cambridge, Massachusetts, USA. This research used resources of the National Energy Research Scientific Computing Center (NERSC). Notebooks producing the plots in this paper can be found at https://github.com/anaismoller/Rubin_KNe_with_Fink_LVK. This research was supported by grants from the Australian Research Council (ARC). AM is supported by the ARC Discovery Early Career Research Award (DE230100055). $S$ is supported by an ARC Discovery Early Career Research Award (DE220100241). Parts of this research were conducted by the ARC Centre of Excellence for Gravitational Wave Discovery (OzGrav), through project number CE230100016.